# Compelling ReLU Network Initialization and Training to Leverage Exponential Scaling with Depth

### v2

@@ Abstract
A neural network with ReLU activations may be viewed as a composition of piecewise linear functions. For such networks, the number of distinct linear regions expressed over the input domain has the potential to scale exponentially with depth, but it is not expected to do so when the initial parameters are chosen randomly. This poor scaling can necessitate the use of overly large models to approximate even simple functions. To address this issue, we introduce a novel training strategy: we first reparameterize the network weights in a manner that forces an exponential number of activation patterns to manifest. Training first on these new parameters provides an initial solution that can later be refined by updating the underlying model weights. This approach allows us to produce function approximations that are several orders of magnitude better than their randomly initialized counterparts.
Machine Learning, ICML
@@ 1 Introduction
Beyond complementary advances in areas like hardware, storage, and networking, the success of neural networks is primarily due to their ability to efficiently capture and represent nonlinear functions (Gibou et al., 2019). In a neural network, the goal of an activation function is to introduce nonlinearity between the network’s layers so that the network does not simplify to a single linear function. The rectified linear unit (ReLU) has a unique interpretation in this regard. Since it can only deactivate a neuron or apply the identity operation, it transforms a series of matrix multiplications into a multi-linear function; each possible configuration of active and inactive neurons can produce a unique linear transformation over a region of input space. Thus, for a ReLU network, the number of activation patterns and their corresponding linear regions can theoretically scale exponentially with the depth of the network [^1] The reader is referred to Chmielewski-Anders (2020) for definitions of linear regions, activation patterns, and activation regions.. Hence deep architectures may outperform shallow ones.
@@
Surprisingly, though, a sophisticated theory of how to best encode functions into ReLU networks is lacking, and in practice, adding depth is often observed to help less than one might expect from this exponential intuition. Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions. Hanin & Rolnick (2019) show a rather disappointing bound pertaining to randomly initialized networks: they prove that the average number of linear regions formed upon initialization is entirely independent of the configuration of the neurons, so depth is not properly utilized. They observed that gradient descent has a difficult time creating new activation regions and that their bounds approximately held after training. As we will discuss later, the number of linear regions is not actually a model property that gradient descent can directly optimize. Gradient descent is also prone to redundancy; Frankle & Carbin (2019) show how up to 90% of neurons may ultimately be eliminated from a network without significantly degrading accuracy.
@@
The present work aims to begin eliminating these inefficiencies, starting in a simple one-dimensional setting. Drawing inspiration from existing theoretical ReLU constructions, our novel contributions include a special reparameterization of a ReLU network that forces it to maintain an exponential number of activation patterns over the input domain. We then demonstrate a novel pretraining strategy, which uses these derived parameters before manipulating the underlying matrix weights. This allows the network to discover solutions that are more accurate and unlikely to be found otherwise. Initializing with an exponential number of regions is already improbable, but our pretraining strategy forcibly retains these regions where unassisted gradient descent would otherwise make short-sighted optimizations that discard them. Thus, our pretraining strategy leaves minimal work up to unassisted gradient descent, which then does not have to discover exponentially many new regions. Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally trained network of equal size. Although the preliminary results in this theoretical study pertain to one-dimensional convex functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
@@ 2 Related Work
This work is primarily concerned with novel training methodology, but it also possesses a significant approximation theoretic component. Our work can be viewed as a first step towards generalizing the approximation to $x^{2}$ we review in this section. The reparameterization we employ modifies this method to become trainable to represent other convex one-dimensional functions and then converts that result back into a matrix representation.
@@ 2.1 Related Work in Approximation
Infinitely wide neural networks are known to be universal function approximators, even with only one hidden layer (Hornik et al., 1989; Cybenko, 1989). Infinitely deep networks of fixed width are universal approximators as well (Lu et al., 2017; Hanin, 2019). In finite cases, the trade-off between width and depth is often studied in assessing the capability of a network to approximate a function.
@@
Notably, there exist functions that can be represented with a sub-exponential number of neurons in a deep architecture, yet which require an exponential number of neurons in a wide and shallow architecture. For example, Telgarsky (2015) shows that deep neural networks with ReLU activations on a one-dimensional input are able to generate symmetric triangle waves with an exponential number of linear segments (shown in Figure 1 as the ReLU network $T(x)$). This network functions as follows: each layer takes a one-dimensional input on $[0,1]$, and outputs a one-dimensional signal also on $[0,1]$. The function they produce in isolation is a single symmetric triangle. Together in a network, each layer feeds its output to the next, performing function composition. Since each layer converts lines from 0 to 1 into triangles, it doubles the number of linear segments in its input signal, exponentially scaling with depth.
@@
The novel reparameterization strategy contributed by the present work aims to extend this idea, compelling ReLU networks to compose rapidly oscillating triangular signals. The key insight is that instead of expressing each layer by its weights, we can train the location on $(0,1)$ of a triangular function’s peak and then set that layer’s weights accordingly. Composing the layers together will create a variety of dilated triangular waveforms that the network can use in its approximations. We emphasize that while the (Telgarsky, 2015) approximation uses fixed symmetric triangles and is not trainable, a key insight of our work is to instead leverage non-symmetric triangles, which yields trainable parameters (one per triangle).
@@
Nonetheless, with only symmetric triangle waves, Yarotsky (2017) and Liang & Srikant (2016) demonstrate the ability to construct $y=x^{2}$ on $[0,1]$ with exponential accuracy. To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1. As these successive approximations are computed, Figure 1 plots their convergence to $x^{2}$, as well as the convergence of the derivatives. We note that these prior papers explicitly focus on approximating $x^{2}$ using symmetric triangle waves, unlike the present work, which is more general (although those papers leverage their $x^{2}$ approximation to also indirectly build approximations of other functions).
@@
The techniques of Yarotsky (2017) and Liang & Srikant (2016) seem to touch on something fundamental regarding how ReLU networks can approximate functions: even though individual neuron outputs may be jagged (e.g., Figure 1, top row), an appropriate infinite sum of them (e.g., $f_{\infty}(x)$) may still be differentiable. Moreover, the derivative of the approximation produced by this technique also converges to the true solution $y=2x$ in the limit. Therefore, inspired by the exponential accuracy of this technique, we place a particular emphasis on investigating the relative importance of differentiability throughout the paper. We will argue with numerical evidence that enforcing convergence of the derivative can be a good protection against overfitting in ReLU networks.
@@
Lastly, other works focus on showing how ReLU networks can encode and subsequently surpass traditional approximation methods (Lu et al., 2021; Daubechies et al., 2022). Interestingly, certain fundamental themes from above like composition, triangles, or squaring are still present. One other interesting comparison of the present work is to Ivanova & Kubat (1995), which uses decision trees as a means to initialize neural networks. It is a sigmoid/classification analogy to this work, but rather than an attempting to improve neural networks with decision trees, it is an attempt to improve decision trees with neural networks.
@@ 2.2 Related Work in Initialization
Our work seeks to improve network initialization by making use of explicit theoretical constructs. This stands in sharp contrast the current standard approach, which treats neurons homogeneously. Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot & Bengio, 2010). They use weight values that are sampled from distributions defined by the input and output dimension of each layer. Aside from sub-optimal approximation power associated with random weights, a common issue is that an entire ReLU network can collapse into outputting a constant value at initialization. This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023). It occurs when the initial weights and biases cause every neuron in a particular layer to output a negative value. The ReLU activation then sets the output of that layer to 0, blocking any gradient updates. Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020). Several papers propose solutions: Shin & Karniadakis (2020) use a data-dependent initialization, while Singh & Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed. We observed during our experiments that RAAI greatly reduces, but does not eliminate the likelihood of dying ReLU. Our approach enforces a specific network structure that does not collapse in this manner.
@@ 3 Compositional Networks
We begin by discussing how to deliberately architect the weights of a 4 neuron wide, depth $d$ ReLU network to induce a number of linear segments exponential in $d$. Throughout the paper, we refer to these as compositional networks. First we will introduce their mathematical form, and then we will discuss how to encode them as ReLU networks.
@@
where $0\leq a_{i}\leq 1$. This produces a triangular shape with a peak at $x=a_{i}$ and both endpoints at $y=0$. Each layer in a compositional network in isolation would compute these if directly fed the input signal. Its derivatives are the piecewise linear functions:
@@
The output of a compositional network will be a weighted sum over the triangle waves formed at each layer. Assuming the network to be infinitely deep we have
@@
Figure 2: The bottom right is a network representation of a triangle function. The bottom left shows that triangle function as a hidden layer of a compositional network. The one-dimensional input and output of a triangle function is instead converted into shared weights within the layers. A full compositional network is assembled above.
To encode these functions into a ReLU network, we begin with triangle functions. This is the subnetwork on the right in Figure 2. Its maximum output is $1$ at the peak location $a\in(0,1)$. Neuron $t_{1}$ simply preserves the input signal. Meanwhile, $t_{2}$ is negatively biased, deactivating it for inputs less than $a$. Subtracting $t_{2}$ from $t_{1}$ changes the slope at the point where $t_{2}$ begins outputting a nonzero signal. The weight $-1/(a-a^{2})=-(1/a+1/(1-a))$ is picked to completely negate $t_{1}$’s positive influence, and then produce a negative slope. When these components are stacked the individual triangles they form will be composed, so it can be beneficial to think about each neuron in terms of its output waveform over the entire input domain $[0,1]$. $t_{1}$ neurons are always active, so they hold complete triangle waves. $t_{2}$ neurons are deactivated for small inputs, so they hold an alternating sequence of triangles and inactive regions. This can be seen in Figure 3. Note how $t_{1}$ and $t_{2}$ use identical slopes where both are active.
@@
Naively stacking the triangle generators from Figure 2 together would form a $1\times 2\times 1\times 2\times 1$ shape, but this is unnecessarily deep. We can replicate the one-dimensional function composition in the hidden layers on the left side of Figure 2 by using weight sharing instead. Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave does so by combining $t_{1}$ and $t_{2}$ in the same proportion. In this way, we can avoid having to use the extra intermediate neurons.
@@
There are two other neurons in the compositional network’s layers. The accumulation neuron passes a weighted sum of all previous triangle waves through each layer. If this were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients. Based on the derivations in Section 3.1, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues. Instead, the ratios between the coefficients are distributed amongst all the weights, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer. A conventional bias will have no connections to prior layers, so it will be unable to adjust to the weight decay. Therefore, a fourth neuron is configured to output a constant signal. Other neurons can then use their connection to it as a bias. This neuron will then connect to itself so that it can scale down with each layer.
3.1 Differentiability of Model Output
Because the reparameterization makes use of exponentially many linear regions, the functions it produces will have high complexity and be in danger of overfitting. To mitigate this we focus on constraining the network to have a differentiable output. While this is technically impossible for a finite network since the output will be piecewise linear; it is possible for a derivative to exist in the limit of infinite depth. We show that there is a unique choice of scaling parameters that is determined by the peaks of the composed triangle functions which can accomplish this. Enforcing a smooth solution during pretraining imparts a bias towards smoothness during the second training stage that will help the model avoid overfit solutions. The condition we derive is necessary for the existence of a derivative. In the appendix we show that it is also sufficient.
Recalling the earlier definition of the network output $F(x)$, we would like to select the $s_{i}$ based on $a_{i}$ in a manner where the derivative $F^{\prime}(x)$ is defined on all of $[0,1]$, which can only be done if the left and right hand derivative limits $F^{\prime}_{+}(x)$ and $F^{\prime}_{-}(x)$ agree.
Notationally, we will denote the sorted $x$-locations of the peaks and valleys of $W_{i}(x)$ by the lists $P_{i}=\{x:W_{i}(x)=1\}$ and $V_{i}=\{x:W_{i}(x)=0\}$. We will use the list $B_{i}$ to reference the locations of all non-differentiable points, which we refer to as bends. $B_{i}:=P_{i}\cup V_{i}$. $f_{i}(x)=\sum_{n=0}^{i-1}s_{n}W_{n}(x)$ will denote finite depth approximations up to but not including layer $i$. The error function $E_{i}(x)=\sum_{n=i}^{\infty}s_{n}W_{n}(x)=F(x)-f_{i}(x)$ will represent the error between the finite approximation and the infinite depth network.
Figure 4: Triangle functions $T_{0}$ and $T_{1}$, and the triangle wave resulting from their composition. Note how $T_{1}$ is reflected in $W_{1}$.
Figure 4 highlights some important properties about composing triangle functions. Peaks alternate with valleys. Peak locations in one layer become valleys in the next, and valleys persist. To produce $W_{i}$, each line segment of $W_{i-1}$ becomes a dilated copy of $T_{i}$. On negative slopes, the input to layer $i$ is reversed, so those copies are reflected. Each triangle function has two distinct slopes $1/a_{i}$ and $-1/(1-a_{i})$ which are dilated by the chain rule during the composition. Because the triangles are situated back to back, the slopes of $W_{i}$ on each side of a valley are proportional. New valleys are scaled by $1/(1-a_{i})$, while old ones are scaled by $1/a_{i}$.
Before we begin reasoning about $F^{\prime}(x)$ we note that it can simplify the analysis to only consider the derivative of the error function $E^{\prime}(x)$.
Lemma 3.1.
for $x\in P_{i}$, $F^{\prime}(x)$ is continuous iff $E^{\prime}_{i}(x)$ is continuous.
Proof.
All of $W_{n}(x)$ for $n<i$ are differentiable at $x\in P_{i}$ since $x$ will lie in the interior of a linear region of $W_{n}$. Therefore, $f^{\prime}_{i}(x)=\sum_{n=0}^{i-1}s_{n}W^{\prime}_{n}(x)$ is continuous. Since $E^{\prime}_{i}+f^{\prime}_{i}=F^{\prime}$, their difference is a continuous function, $F^{\prime}(x)$ is continuous iff $E^{\prime}_{i}(x)$ is continuous. ∎
Here we compute the right derivative $(E_{i})^{\prime}_{+}(x)$ of the error function. The left derivative will only be different by a constant factor.
Lemma 3.2.
For all $x\in P_{i}$, $E_{+}^{\prime}(x)$ and $E_{-}^{\prime}(x)$ are proportional to
$s_{i}-\frac{1}{1-a_{i+1}}\left(s_{i+1}+\sum_{n=i+2}^{\infty}s_{n}\prod_{k=i+2}^{n}\frac{1}{a_{k}}\right).$ (4)
Proof.
Let $x_{n}$ be some point in $P_{i}$, and let $n$ be its index in any list it appears in. To calculate the value of $E_{+}^{\prime}(x)=\sum_{n=i}^{\infty}s_{n}(W_{n})^{\prime}_{+}(x)$, we will have to find the slope of the linear intervals to the immediate right of $x_{n}$ for all $W_{i}$. The terms in the summation (4) mostly derive from the chain rule. We will use $R$ to represent $W_{i+}^{\prime}(x)$. The first term in the sum will be $Rs_{i}$. $W^{\prime}_{i+1}(x)=T^{\prime}_{i+1}(W_{i}(x))W^{\prime}_{i}(x)$ since the derivatives of composed functions will multiply. There are two different possible slope values of $T_{i+1}$ to multiply by, and the correct one to multiply is $-1/(1-a_{i+1})$ because $x_{n}$ is a peak of $W_{i}$, so $W_{i}(x)>a_{i+1}$ for $x\in(B_{i+1}[n-1],B_{i+1}[n+1])$. This gives $W^{\prime}_{i+1}(x)=-R\frac{s_{i+1}}{(1-a_{i+1})}$. Note that the second term has the opposite sign as the first.
For all remaining terms, since $x_{n}$ was in $P_{i}$, it is in $V_{j}$ for $j>i$. For $x\in(B_{j+1}[n-1],B_{j+1}[n+1])$, $W_{j}(x)<a_{j+1}$ and the chain rule applies the first slope $1/a_{j+1}$. Since this slope is positive, every term has the opposite sign as the first. Summing up all the terms with the coefficients $s_{i}$, and factoring out $R$ will yield the desired formula. Note that this same reasoning applies to the left side, the initial slope $L$ will just be different. ∎
Lemma 3.3.
If $E^{\prime}(x)$ exists, it must be equal to 0.
Proof.
Let $S$ represent Equation (4), and $R$ and $L$ be the constants of proportionality for the directional derivatives. If $E^{\prime}_{+}=E^{\prime}_{-}$, then $R_{x}S=L_{x}S$ for all $x\in P_{i}$. Since $W_{i}$ is comprised of alternating positive and negatively sloped line segments, $R_{x}$ and $L_{x}$ have opposite signs. The only way to satisfy the equation then is if $S=0$. Consequently, $E^{\prime}(x)=0$ for all $x\in P_{i}$. ∎
We can now prove our main theorem. Intuitively, the theorem shows that there is a way to sum the triangular waveforms $W_{i}$ so that the resulting function approximation converges to a differentiable function, which, as mentioned at the start of this subsection, can aid in preventing overfitting when using exponentially many linear regions. The idea of the proof is that much of the formula for $E^{\prime}(x)$ will be shared between two successive generations of peaks. Once they are both valleys, they will behave the same, so the sizes of their remaining discontinuities will need to be proportional.

### v3

@@ Abstract
A neural network with ReLU activations may be viewed as a composition of piecewise linear functions. For such networks, the number of distinct linear regions expressed over the input domain has the potential to scale exponentially with depth, but it is not expected to do so when the initial parameters are chosen randomly. This poor scaling can necessitate the use of overly large models to approximate even simple functions. To address this issue, we introduce a novel training strategy: we first reparameterize the network weights in a manner that forces the network to display a number of activation patterns exponential in depth. Training first on our derived parameters provides an initial solution that can later be refined by directly updating the underlying model weights. This approach allows us to learn approximations of convex, one-dimensional functions that are several orders of magnitude more accurate than their randomly initialized counterparts.
@@ 1 Introduction
Beyond complementary advances in areas like hardware, storage, and networking, the success of neural networks is primarily due to their ability to efficiently capture and represent nonlinear functions (Gibou et al., 2019). In a neural network, the goal of an activation function is to introduce nonlinearity between the network’s layers so that the network does not simplify to a single linear function. The rectified linear unit (ReLU) has a unique interpretation in this regard. Since it either deactivates a neuron or acts as an identity, the resulting transformation on each individual input remains linear. However, each possible configuration of active and inactive neurons can produce a unique linear transformation over a particular region of input space. The number of these activation patterns and their corresponding linear regions provides a way to measure the expressivity of a ReLU network, and can theoretically scale exponentially with the depth of the network [^1] The reader is referred to Chmielewski-Anders (2020) for definitions of linear regions, activation patterns, and activation regions.. Hence deep architectures may outperform shallow ones.
@@
Surprisingly, though, a sophisticated theory of how to best encode functions into ReLU networks is lacking, and in practice, adding depth is often observed to help less than one might expect from this exponential intuition. Lacking more advanced theory, practitioners typically use random parameter initialization and gradient descent, the drawbacks of which often lead to extremely inefficient solutions. Hanin and Rolnick (2019) show a rather disappointing bound pertaining to randomly initialized networks: they prove that the average number of linear regions formed upon initialization is entirely independent of the configuration of the neurons, so depth is not properly utilized. They observed that gradient descent has a difficult time creating new activation regions and that their bounds approximately held after training. As we will discuss later, the number of linear regions is not actually a model property that gradient descent can directly optimize. Gradient descent is also prone to redundancy; Frankle and Carbin (2019) show how up to 90% of neurons may ultimately be eliminated from a network without significantly degrading accuracy.
@@
The present work aims to begin eliminating these inefficiencies, starting in a simple one-dimensional setting. Drawing inspiration from existing theoretical ReLU constructions, our novel contributions include a special reparameterization of a ReLU network that forces it to maintain an exponential number of activation patterns over the input domain. We then demonstrate a novel pretraining strategy, which trains these derived parameters before manipulating the underlying matrix weights. This allows the network to discover solutions that are more accurate and unlikely to be found otherwise. Our technique directly rectifies the issues raised by Hanin and Rolnick (2019). Firstly, we initialize immediately with an exponential number of linear regions, which would not be expected to happen otherwise. Secondly, we are not reliant on gradient descent to “discover” new regions; we only need to maintain the existing ones, which is already guaranteed by the reparameterization during the pretraining stage. Additionally, this pretraining step can avoid areas of the loss landscape where unassisted gradient descent might make short-sighted optimizations that eliminate activation regions. Our results demonstrate that minimizing the reliance of network training on unassisted gradient descent can reliably produce error values orders of magnitude lower than a traditionally-trained network of equal size. Although the preliminary results in this theoretical study pertain to one-dimensional convex functions, the paper concludes with our views on extending these theoretical exponential benefits for ReLU networks to arbitrary smooth functions with arbitrary dimensionality, which would have significant practical utility.
@@ 2 Related Work
This work is primarily concerned with a novel training methodology, but it also possesses a significant approximation theoretic component. Our work can be viewed as a first step towards generalizing the approximation to $x^{2}$ we review in this section. The reparameterization we employ modifies this method to become trainable in order to represent other convex one-dimensional functions, and then converts that result back into a matrix representation.
@@ 2.1 Related Work in Approximation
Infinitely wide neural networks are known to be universal function approximators, even with only one hidden layer (Hornik et al., 1989; Cybenko, 1989). Infinitely deep networks of fixed width are universal approximators as well (Lu et al., 2017; Hanin, 2019). In finite cases, the trade-off between width and depth is often studied in assessing a network’s ability to approximate (learn) a function.
@@
Notably, there exist functions that can be represented with a sub-exponential number of neurons in a deep architecture, yet which require an exponential number of neurons in a wide and shallow architecture. For example, Telgarsky (2015) shows that deep neural networks with ReLU activations on a one-dimensional input are able to generate symmetric triangle waves with an exponential number of linear segments (shown in Figure 1 as the ReLU network $T(x)$). This network functions as follows: each layer takes a one-dimensional input on $[0,1]$, and outputs a one-dimensional signal also on $[0,1]$. The function they produce in isolation is a single symmetric triangle. Together in a network, each layer inputs its output to the next, performing function composition. Since each layer converts lines from 0 to 1 into triangles, it doubles the number of linear segments in its input signal, exponentially scaling with depth.
@@
The novel reparameterization strategy contributed by the present work aims to extend this idea, compelling ReLU networks to compose rapidly oscillating triangular signals. The key insight is that instead of expressing each layer by its weights, we can train the location on $(0,1)$ of a triangular function’s peak and then set that layer’s weights accordingly. Composing the layers together will create a variety of dilated triangular waveforms that the network can use in its approximations. We emphasize that while the Telgarsky (2015) approximation uses fixed symmetric triangles and is not trainable, a key insight of our work is to instead leverage non-symmetric triangles, which yields trainable parameters (one per triangle).
@@
Nonetheless, with only symmetric triangle waves, Yarotsky (2017) and Liang and Srikant (2016) demonstrate the ability to construct $y=x^{2}$ on $[0,1]$ with exponential accuracy. To produce their approximation, one begins with $f_{0}(x)=x$, then computes $f_{1}(x)=f_{0}(x)-T(x)/4$, $f_{2}(x)=f_{1}(x)-T(T(x))/16$, $f_{3}(x)=f_{2}(x)-T(T(T(x)))/64$, and so forth, as pictured in Figure 1. As these successive approximations are computed, Figure 1 plots their convergence to $x^{2}$, as well as the convergence of the derivatives. We note that these prior papers explicitly focus on approximating $x^{2}$ using symmetric triangle waves, unlike the present work, which is more general (although those papers leverage their $x^{2}$ approximation to also indirectly build approximations of other functions).
@@
The techniques of Yarotsky (2017) and Liang and Srikant (2016) seem to touch on something fundamental regarding how ReLU networks can approximate functions: even though individual neuron outputs may be jagged (e.g., Figure 1, top row), an appropriate infinite sum of them (e.g., $f_{\infty}(x)$) may still be differentiable. Moreover, the derivative of the approximation produced by this technique also converges to the true solution $y=2x$ in the limit. Therefore, inspired by the exponential accuracy of this technique, we place a particular emphasis on investigating the relative importance of differentiability throughout the paper. We will argue with numerical evidence that enforcing convergence of the derivative can reduce the number of local minima during training.
@@
Lastly, other works focus on showing how ReLU networks can encode and subsequently surpass traditional approximation methods (Lu et al., 2021; Daubechies et al., 2022). Interestingly, certain fundamental themes from above like composition, triangles, or squaring are still present. One other interesting comparison of the present work is to Ivanova and Kubat (1995), which uses decision trees as a means to initialize neural networks. It is a sigmoid/classification analogy to this work, but rather than an attempting to improve neural networks with decision trees, it is an attempt to improve decision trees with neural networks.
@@ 2.2 Related Work in Initialization
Our work seeks to improve network initialization by making use of explicit theoretical constructs. This stands in sharp contrast the current standard approach, which treats neurons homogeneously. Two popular initialization methods implemented in PyTorch are the Kaiming (He et al., 2015) and Xavier initialization (Glorot and Bengio, 2010). They use weight values that are sampled from distributions defined by the input and output dimension of each layer. Aside from sub-optimal approximation power associated with random weights, a common issue is that an entire ReLU network can collapse into outputting a constant value at initialization. This is referred to as the dying ReLU phenomenon (Qi et al., 2024; Nag et al., 2023). It occurs when the initial weights and biases cause every neuron in a particular layer to output a negative value. The ReLU activation then sets the output of that layer to 0, blocking any gradient updates. Worryingly, as depth goes to infinity, the dying ReLU phenomenon becomes increasingly likely (Lu et al., 2020). Several papers propose solutions: Shin and Karniadakis (2020) use a data-dependent initialization, while Singh and Sreejith (2021) introduce an alternate weight distribution called RAAI that can reduce the likelihood of the issue and increase training speed. We observed during our experiments that RAAI greatly reduces, but does not eliminate the likelihood of dying ReLU. Our approach enforces a specific network structure that does not collapse in this manner.
@@ 3 Compositional Networks
We begin by discussing how to deliberately architect the weights of a 4-neuron-wide, depth $d$ ReLU network to induce a number of linear segments exponential in $d$. Throughout the paper, we refer to these as compositional networks. First we will introduce their mathematical form, and then we will discuss how to encode them as ReLU networks.
@@
where $0\leq a_{i}\leq 1$. This produces a triangular shape with a peak at $x=a_{i}$ and with both endpoints at $y=0$. Each layer in a compositional network in isolation would compute these if directly fed the input signal. $T_{i}(x)$’s derivatives are the piecewise linear functions:
@@
The output of a compositional network will be a weighted sum over the triangle waves formed at each layer. Assuming the network to be infinitely deep, we have
@@
Figure 2: The rightmost diagram is a network representation of a triangle function. The middle shows that triangle function as a hidden layer of a compositional network. The one-dimensional input and output of a triangle function is instead converted into shared weights within the layers. A full compositional network is assembled on the left.
To encode these functions into a ReLU network, we begin with triangle functions. This is the subnetwork on the right in Figure 2. Its maximum output is $1$ at the peak location $a\in(0,1)$. Neuron $t_{1}$ simply preserves the input signal. Meanwhile, $t_{2}$ is negatively biased, deactivating it for inputs less than $a$. Subtracting $t_{2}$ from $t_{1}$ changes the slope at the point where $t_{2}$ begins outputting a nonzero signal. The weight $-1/(a-a^{2})=-(1/a+1/(1-a))$ is picked to completely negate $t_{1}$’s positive influence, and then produce a negative slope. When these components are combined into a deep network, the individual triangles they form will be composed with those earlier in the network, but the two neurons will still behave analogously. Considering the output waveform of $t_{1}$ and $t_{2}$ neurons over the entire input domain $[0,1]$ and at arbitrary depth, we have that $t_{1}$ neurons are always active, outputting complete triangle waves. In contrast, $t_{2}$ neurons are deactivated for small inputs, so they output an alternating sequence of triangles and inactive regions, which can be seen in Figure 3.
@@
Naively using the output of one triangle generator (from Figure 2) as input to the next would form a $1\times 2\times 1\times 2\times 1$ shape, but this is unnecessarily deep. We can replicate the one-dimensional function composition in the hidden layers on the left side of Figure 2 by using weight sharing instead. Any outgoing weight from $t_{1}$ or $t_{2}$ is shared; every neuron taking in a triangle wave as input does so by combining $t_{1}$ and $t_{2}$ in the same proportion. In this way, we can avoid having to use the extra intermediate neurons. Notice in Figure 3 how the waves output by $t_{1}$ and $t_{2}$ neurons always have identical slopes in regions where both neurons are active.
@@
There are two other neurons in the compositional network’s layers. The accumulation neuron (marked as “sum” in Figure 2) acts similarly to a skip connection, maintaining a weighted sum of all previous triangle waves through each layer. If this were naively implemented, it would multiply the $t_{1}$ and $t_{2}$ weights by the sum coefficients. Based on the derivations in the appendix, these coefficients will be exponentially decaying, so learning these weights directly may cause conditioning issues. Instead, all weights in each layer are multiplied by the ratios between successive scaling coefficients, so that the outputs of $t_{1}$ and $t_{2}$ neurons decay in amplitude in each layer. A conventional bias will have no connections to prior layers, so it will be unable to adjust to the rescaling neuron outputs without having to learn exponentially small values. Therefore, a fourth neuron is configured to output a constant signal, so the weight other neurons place on this output can replace their bias. The bias neuron will connect to the previous layer’s bias neuron so that the constant signal can scale down gradually with each layer.
3.1 Differentiable Model Output
Given the rapid oscillations of the triangle waves formed at each layer, the resulting output of the network will appear jagged with most choices of scaling parameters. This can lead to the network settling into overfit local minima, in which the data points might lie along the curve, but the curve is not predictive of the areas in between due to rapid oscillation. Our main mathematical result addresses this issue by forcing differentiability of the network output. It turns out that for this class of functions, the peak locations of the triangle waves uniquely determine the scales with which to sum them.